# RealityCraft: An In-Situ CAD+CAM Interface for Novices via Scene-Aware Augmented Reality

### v1

@@ 4.2.5. Solid and Hole Objects
Users can toggle whether an object is solid, adding material, or hole, removing material, by pressing a button on their left wrist. The hole objects appear with a transparent mask, making them identifiable when compared to the solid objects (Figure 6a). The hole objects can be manipulated the same way as the solid objects.
@@ 4.2.6. Combining Objects
Users can combine and turn selected objects into one uniform object with a button on the left wrist. The combined object will turn into the color of the last solid object chosen. The combining operation works with solid and hole objects, removing the hole objects after performing a cut on the solid objects through the intersection between the solid and hole objects (Figure 6b and 6c).
@@ 4.2.7. Ruler Tool
Users can create a virtual ruler by pressing a button on their left wrist. This simple ruler, equipped with two interactable edge handles (similar to scaling handles), provides real-world scale measurements. This ruler can be used for precise alignment and measuring distances between objects within the real and virtual space (Figure 8a).
@@ 4.2.8. Parametric Resizing
In addition to free, snap-to-grid, and uniform scaling operations, users can also adjust object sizes through precise numerical length inputs. To use this feature, users first select the edge of the Object Manipulation Box they wish to resize (Figure 8b) using a pinch gesture. Then, a numerical keyboard appears on the user’s left hand (Figure 8c), allowing them to enter the desired real-world length for the selected edge. The object scales proportionally along the direction of the chosen edge based on this input (Figure 8d). This method can be effectively combined with the ruler tool to ensure accurate design according to specific measurements.
@@ 4.2.9. Other Functions
Other buttons placed on the right wrist enhance the efficiency of the design process. These features include selecting the color and material of the objects, duplicating selected objects, selecting or deselecting all objects, deleting selected objects, deleting all objects, undo, redo, saving, and loading operations.
@@ 4.3. Fabrication Features
The implementation of the server system was possible with the usage of Docker containers to package the working environment for different computers. The server itself was written with JavaScript in Node.js, and the slicer used for the server was Ultimaker Cura111111https://ultimaker.com/software/ultimaker-cura/.
Four images labeled (a), (b), (c), and (d) show a user’s first-person view looking on top of a table where there is a horizontal grid appearing on the table and a virtual 3D printer twin and a model of a wall hook is visible. The 3D printer twin is seen as a transparent rectangular prism with white edges and a grid on the bottom plate. The name of the printer (Creality Ender-3 V3 KE) is written on the bottom of the printer. In the picture (a), the user is moving the model with the movement handle grabbed with a pinch gesture with the right hand, dragging it to the printer twin. A white line grid appears on the printer twin. In picture (b), the model is placed in the middle of the printer twin and it says ”model printed successfully” in green color on top of the printer twin. On the right side of the printer twin, two input fields and two buttons are visible. The first input field is labeled ”enter sever IP…” and the second input field is labeled ”enter printer IP…”. The first button is labeled ”Start print” and the second button is labeled ”Slicer settings”. In the picture (c), user is rotating the model with a pinch gesture applied to the rotation handle on top of the model. In picture (d), a progress bar is seen on top of the printer twin and a text is written as ”0%” in the middle. The input fields contain the partially blurred IP addresses of the server and the printer and the buttons are labeled ”pause print” and ”abort print”.
@@ 4.3.3. Preparing for Printing and Slicing
A client can make multiple requests on the server. When a /slice POST request is sent to the server, the STL file is loaded on the appropriate slicer, and the G-code file of the target model is obtained. After that, the printing process can start. Note that one server can handle multiple printers, and one headset can connect to various servers. The only necessary action is to enter IP addresses.
@@ 4.3.4. Printing and Monitoring
After the /print POST request is sent to the virtual printer twin object, the server will connect and get access permission to the 3D printer. This permission is only necessary the first time as the server will save the token for accessing the printer, and access permission will not be necessary until the token has expired. If access to the 3D printer is granted, the server will send a printing request with the G-code file to the printer. The headset can obtain the print process information by sending the /print GET request. The user can pause or abort the printing process by sending a PUT request to the /print URL if necessary. This request supports three different commands: continue, pause, and stop, which can be sent with buttons next to the printer twin (Figure 9d).
@@ 5. Applications
In this section, we present ten distinct designs created with RealityCraft (Figure 10) to demonstrate its versatility and capabilities across various workspace environments. These examples highlight how RealityCraft enables users to easily produce both practical and creative objects through spatial interaction, tailored to their immediate surroundings.
@@ 5.1. Desk Workspace
In the desk workspace, five designs were developed: a phone stand, drink coaster, pen holder, candy bowl, and fidget terrain (Figure 10a-e). These objects demonstrate RealityCraft’s ability to facilitate ergonomic, personalized designs suited for everyday use. The phone stand and drink coaster highlight basic cutting operations and represent simple designs in everyday objects. The pen holder showcases intricate decorative patterns, made by combining and duplicating hole objects in a chain, and the candy bowl explores experimental shapes by combining rectangular prisms into curvilinear forms. Fidget terrain stands out as a playful interactive object designed with ramps and rails, which illustrates our system’s potential for creating designs that interact with moving physical objects. Each design benefits from real-time in-situ customization, ensuring the objects fit and function precisely within the user’s workspace. Here, working on smaller scales with hand interactions can be challenging; thus, a useful approach is to create the design on a larger scale and then scale it down to the intended size. This method helps manage precision and detail when working with smaller components.
@@ 5.2. Wall Workspace
For the wall workspace, two objects were designed: a figurine stand and a key hanger (Figure 10f and 10g). The figurine stand allows users to display small decorative items without adding clutter, created using the ruler tool for a fitting hole. The key hanger offers practical functionality with aligned hooks using snap-to-grid. The models are snapped to the wall while designing for a better sense of depth, helping users create visually integrated solutions.
@@ 5.3. Other Workspaces
Beyond the more common desk and wall spaces, RealityCraft can also be applied to other environments, e.g., ground, ceiling, doors, windows, screens, couches, and more. On the ground, a door wedge was designed to hold doors securely in place (Figure 10h), benefiting from real-time scaling for accurate size.
For the ceiling, we created two distinct designs: a plant pot hanger and a light shade (Figure 10i and 10j). The plant pot hanger demonstrates the system’s ability to accommodate objects that require support and suspension, such as using ”ropes” for hanging, a practical solution for overhead applications. The light shade, on the other hand, highlights the system’s capability to handle complex and abstract 3D modeling. Its artistic design illustrates the system’s flexibility in creating intricate structures on a headset without technical limitations. These examples show the adaptability to less conventional design spaces and orientations.
@@ 6. Evaluation
To evaluate RealityCraft, we conducted a user study involving three design tasks of increasing complexity. We used both quantitative measures and qualitative feedback to gauge user performance.
@@ 6.1. Methods
Participants. We recruited seven participants (ages 21-25) with varying levels of experience in CAD and AR systems. Three participants had little experience with CAD software (e.g., Tinkercad), and four had no previous CAD experience. Additionally, two participants had no prior experience with AR systems, four had used AR at some point, and one worked as an AR developer in the past.
Two images labeled (a) and (b) show twelve objects each: four pen holders, four finger skateboard ramps, and four wall-mounted key hangers with three hooks on each. On the image (a), the digital renders of the models can be seen. On the image (b), the 3D-printed results are seen on top of a desk and the wall.
Task & Procedure. For the evaluation procedure, we assigned participants modeling tasks with specific requests, and after they finished their tasks and printed their models, we conducted a survey. We assigned three modeling tasks to each participant. The first task was designing a pen holder, i.e., a relatively easy crafting task to accustom users to the interface. Here, we requested users to ensure the pens can fit into the holder without falling. The second task was designing a fidget toy ramp obstacle where users need to integrate their designs with moving real-world objects. Users are requested to design with a height limit of 10 cm. The third task was designing a wall-mounted key hanger with three hooks, which are requested to be narrow enough for keychain rings to go through. This task carries the challenge of designing with smaller details. Participants used a Quest 3 headset and interacted with the system through hand gestures. Participants started their prints on the virtual printer twins. For the sake of time, we did not make them wait until the printer finished but proceeded to the next tasks.
@@
Measures. We recorded the time and the amount of CSG boolean operations (combine operations) required to complete the individual tasks as a measure of overall system performance. After completing the tasks, participants were asked to complete standardized questionnaires, including the System Usability Scale (SUS) (Brooke, 1996), which we extended with additional questions tailored to our system, the short version of the User Experience Questionnaire (UEQ) (Laugwitz et al., 2008), and the NASA Task Load Index (NASA-TLX) (Hart and Staveland, 1988). Additionally, we collected qualitative feedback through open-ended questions regarding their overall experience.
@@ 6.2. Results
6.2.1. Time and Number of CSG Boolean Operations
Across all tasks, the average total completion time per user was 15m 44s (SD=5m 52s) and the average number of boolean operations was 12.7 (SD=6.7). The first task, crafting a pen holder, had an average time of 3m 5s (SD=1m 22s) and an average number of boolean operations of 2 (SD=1.4). The second task, the fidget toy obstacle, took an average of 5m 53s (SD=2m 44s) and the average number of boolean operations was 3.9 (SD=2.5). Finally, the wall-mounted key holder took 6m 46s on average (SD=2m 8s) with an average number of boolean operations of 6.9 (SD=4). The increasing task completion times and the amount of CSG boolean operations reflect the growing complexity of the designs.
@@ 6.2.2. Observations and Analysis of the Fabricated Artifacts
Figure 11 shows the three types of objects designed and fabricated by the participants. All of the users were successful with the task requirements. Some notable insights were that cube primitive as a solid and cylinder geometry as a hole were mostly used objects. Snap-to-grid was heavily utilized, especially on the key hanger task. All of the participants used the ruler tool in the fidget toy ramp task. Participants relied heavily on depth occlusion and the highlighted projection of the objects on the grid to understand the location and depth of the objects. Seen in Figure 11a, the most complex outcomes were the pen holder by P4, the ramp by P6, and the key hanger by P3. For the pen holder, P4 cut the top part of a pyramid geometry, turned it upside down, and cut a hole inside. For the ramp, P6 cut a cube from multiple angles, finishing in 3m 31s and with 8 boolean operations. For the key hanger, P3 turned a cylinder into a thin disk, cut it from different sides to get a hook shape, and duplicated this hook to put it on a thin plate, aligning the hooks with the snap-to-grid feature. This task took 11m 10s and was done with 15 CSG boolean operations.

### v2

@@ 4.2.6. Solid and Hole Objects
Users can toggle whether an object is solid, adding material, or hole, removing material, by pressing a button on their weak hand wrist. The hole objects appear with a transparent mask, making them identifiable when compared to the solid objects (Figure 9a). The hole objects can be manipulated the same way as the solid objects.
@@ 4.2.7. Combining Objects
Users can combine and turn selected objects into one uniform object with a button on the weak hand wrist. The combined object will turn into the color of the last solid object chosen. The combining operation works with solid and hole objects, removing the hole objects after performing a cut on the solid objects through the intersection between the solid and hole objects (Figure 9b and 9c).
Three photos labeled (a), (b), and (c) show a first-person view of a user looking at a wall. In the image (a) a solid cube and a hole sphere intersecting with the top corner of the cube is seen. The hole (or negative) sphere is seen with red color and with dotted transparent skin. In image (b) the weak hand wrist is visible while the cube and the sphere are selected and colored cyan. The user is about to poke a button named ”combine” with their dominant-hand index finger. In the image (c), the combination of the objects is visible where there is a round cut on the upper corner of the cube where the sphere used to intersect the cube.
4.2.8. Reference Objects
TinkerXR supports creating reference objects, offering users a powerful tool for designing around physical items. Users can activate the reference mode via the hand menu on their dominant hand wrist (Figure 10a), which converts default block objects into reference objects. These reference objects are transparent and feature highlighted wireframe edges, distinguishing them from standard objects.
This wireframe structure allows users to precisely position grids on reference objects, aligning them with vertices and plane normals for accurate design. This translates to the grid not only functioning horizontally or vertically (as seen on the common workspaces) but also diagonally in any direction, fitting to the selected surface of the reference object. The transparency aids in manipulating these objects to replicate physical counterparts (Figure 10b,c), enabling users to build around them easily. Additionally, reference objects can be set as hole types, letting users capture intricate details of physical objects for better integration.
@@ 4.2.9. Ruler Tool
Users can create a virtual ruler by pressing a button on their weak hand wrist. This simple ruler, equipped with two interactable edge handles (similar to scaling handles), provides real-world scale measurements. This ruler can be used for precise alignment and measuring distances between objects within the real and virtual space (Figure 11a).
Four images labeled (a), (b), (c), and (d) show a user’s first-person view using the system. In picture (a), user is using their weak and dominant hand to pinch both ends of a digital line where in the middle of the line it says ”20.00 cm”. The user is measuring out the length of a game console (Nintendo Switch). In pictures (b), (c), and (d) user is manipulating a digital cube model. In the picture (b) user is applying a pinch motion to a small black handle on the middle of the edge of the cube. In picture (c), a numerical keyboard interface appears on the weak hand in direction of the palm and it says ”10 cm” in the input field of the keyboard interface. In picture (d), the user has inputted ”20 cm” as the input and clicks the confirmation button. The cube’s width is doubled.
@@ 4.2.10. Parametric Resizing
In addition to free, snap-to-grid, and uniform scaling operations, users can also adjust object sizes through precise numerical length inputs. To use this feature, users first select the edge of the Object Manipulation Box they wish to resize (Figure 11b) using a pinch gesture. Then, a numerical keyboard appears on the user’s weak hand (Figure 11c), allowing them to enter the desired real-world length for the selected edge. The object scales proportionally along the direction of the chosen edge based on this input (Figure 11d). This method can be effectively combined with the ruler tool to ensure accurate design according to specific measurements.
@@ 4.2.11. Other Functions
Other buttons placed on the dominant hand wrist enhance the efficiency of the design process. These features include selecting the color and material of the objects, duplicating selected objects, selecting or deselecting all objects, deleting selected objects, deleting all objects, undo, redo, saving, and loading operations.
@@ 4.3. Fabrication Features
The implementation of the server system was possible with the usage of Docker containers to package the working environment for different computers. The server itself was written with JavaScript in Node.js, and the slicer used for the server was Ultimaker CuraEngine111111https://ultimaker.com/software/ultimaker-cura/.
Four images labeled (a), (b), (c), and (d) show a user’s first-person view looking on top of a table where there is a horizontal grid appearing on the table and a virtual 3D printer twin and a model of a wall hook is visible. The 3D printer twin is seen as a transparent rectangular prism with white edges and a grid on the bottom plate. The name of the printer (Creality Ender-3 V3 KE) is written on the bottom of the printer. In the picture (a), the user is moving the model with the movement handle grabbed with a pinch gesture with the dominant hand, dragging it to the printer twin. A white line grid appears on the printer twin. In picture (b), the model is placed in the middle of the printer twin and it says ”model printed successfully” in green color on top of the printer twin. On the right side of the printer twin, two input fields and two buttons are visible. The first input field is labeled ”enter sever IP…” and the second input field is labeled ”enter printer IP…”. The first button is labeled ”Start print” and the second button is labeled ”Slicer settings”. In the picture (c), user is rotating the model with a pinch gesture applied to the rotation handle on top of the model. In picture (d), a progress bar is seen on top of the printer twin and a text is written as ”0%” in the middle. The input fields contain the partially blurred IP addresses of the server and the printer and the buttons are labeled ”pause print” and ”abort print”.
The low-poly Stanford bunny is placed inside the virtual printer twin with slicing options displayed next to it.
@@ 4.3.3. Preparing for Printing and Slicing
The sliced model is visualized in the virtual printer (Figure 13b and 13c), where the main walls are displayed in gray and support structures in cyan (tree support selected). The visualization allows users to review the slicing results and identify areas requiring adjustment before starting the print. If the selected model exceeds the printer’s build area, the system notifies users by displaying a ”too large for printer” indicator above the virtual printer twin (Figure 13d). This ensures that users can modify their design or scale it appropriately before initiating the print.
@@
The server manages multiple slicers to serve the XR headset clients. An example slicer configuration for testing has been created for Ultimaker S3121212https://ultimaker.com/3d-printers/s-series/ultimaker-s3/ printer with options: Enable Support, Support Structure, Layer Height, Infill Density, and Infill Pattern (Figure 13a). More slicers can be implemented by adding them to the slicer_options.json file. The slicer options are completely modifiable and could be increased with ease as the run command is also read from the JSON. How the slicer should work is completely in the hands of the programmer as the command could be easily manipulated by the flexible replacement tags that have been noted as options. The only required arguments that a slicer should take are output and input file arguments, which use gcode_path and stl_path, respectively.
A client can make multiple requests on the server. When a /slice GET request is sent to the server, it returns the available slicers and their options. On the other hand, when a /slice POST request is sent to the server, the STL file is loaded on the appropriate slicer, and the G-code file of the target model is obtained. The server parses the options selected by the user and passes them to the slicer. After the slicing operation has been completed, the resulting G-code file is visualized in the virtual printer. This draft object is shown in different colors for the main and the support parts. After that, the printing process can start. Note that one server can handle multiple printers, and one headset can connect to various servers. The only necessary action is to enter IP addresses.
@@ 4.3.4. Printing and Monitoring
After the /print POST request is sent to the virtual printer twin object, the server will connect and get access permission to the 3D printer. This permission is only necessary the first time as the server will save the token for accessing the printer, and access permission will not be necessary until the token has expired. If access to the 3D printer is granted, the server will send a printing request with the G-code file to the printer. The headset can obtain the print process information by sending the /print GET request. The user can pause or abort the printing process by sending a PUT request to the /print URL if necessary. This request supports three different commands: continue, pause, and stop, which can be sent via the buttons next to the printer twin (Figure 12d).
@@ 5. Applications
In this section, we present ten distinct designs created with TinkerXR (Figure 14) to demonstrate its versatility and capabilities across various workspace environments. These examples highlight how TinkerXR enables users to easily produce both practical and creative objects through spatial interaction, tailored to their immediate surroundings.
@@ 5.1. Desk Workspace
In the desk workspace, five designs were developed: a phone stand, drink coaster, pen holder, candy bowl, and fidget terrain (Figure 14a-e). These objects demonstrate the system’s ability to facilitate personalized designs suited for everyday use. The phone stand and drink coaster highlight basic cutting operations and represent simple designs in everyday objects. The pen holder showcases intricate decorative patterns, made by combining and duplicating hole objects in a chain, and the candy bowl explores experimental shapes by combining rectangular prisms into curvilinear forms. Fidget terrain stands out as a playful interactive object designed with ramps and rails, which illustrates the potential for creating designs that interact with moving physical objects. Each design benefits from real-time in-situ customization, ensuring the objects fit and function precisely within the user’s workspace. Here, working on smaller scales with hand interactions can be challenging; thus, a useful approach is to create the design on a larger scale and then scale it down to the intended size. This helps manage precision and detail when working with smaller components.
@@ 5.2. Wall Workspace
For the wall workspace, two objects were designed: a figurine stand and a key hanger (Figure 14f and 14g). The figurine stand allows users to display small decorative items without adding clutter, created using the ruler tool for a fitting hole. The key hanger offers practical functionality with aligned hooks using snap-to-grid. The models are snapped to the wall while designing for a better sense of depth, helping users create visually integrated solutions.
@@ 5.3. Other Workspaces
Beyond the more common desk and wall spaces, TinkerXR can also be applied to other environments, e.g., ground, ceiling, doors, windows, screens, couches, and more. On the ground, a door wedge was designed to hold doors securely in place (Figure 14h), benefiting from real-time scaling for accurate size.
For the ceiling, we created two distinct designs: a plant pot hanger and a light shade (Figure 14i and 14j). The plant pot hanger demonstrates the system’s ability to accommodate objects that require support and suspension, such as using ”ropes” for hanging, a practical solution for overhead applications. The light shade, on the other hand, highlights the system’s capability to handle complex and abstract 3D modeling. Its artistic design illustrates the system’s flexibility in creating intricate structures on a headset without technical limitations. These examples show the adaptability to less conventional design spaces and orientations.
@@ 6. Evaluation
To evaluate TinkerXR, we conducted a user study involving three design tasks of varying contexts, comparing to Tinkercad as a baseline. We used both quantitative measures and qualitative feedback to gauge user performance.
@@ 6.1. Methods
Participants. We recruited ten participants (ages 23-29; 9 male, 1 female). Four participants had little experience with CAD software, and six had no previous CAD experience. Additionally, three participants had little experience with AR systems, and seven had no experience with AR systems. Finally, four have used 3D printers in some parts of their life, while six have never used a 3D printer.
Task & Procedure. For the evaluation, we conducted a within-subjects comparative study to assess TinkerXR against Tinkercad. Each participant completed modeling tasks using both tools, with the order of tool usage counterbalanced across participants to mitigate order effects (e.g., half of the participants used TinkerXR first, followed by Tinkercad, while the other half did the opposite). We showed both systems briefly, and participants completed a 3-minute warm-up before starting the tasks to familiarize themselves with the tools. Participants were assigned three modeling tasks, each tailored to a specific context and accompanied by explicit design requirements: #1 Wall-mounted key hanger: Participants created a key hanger with two hooks narrow enough to accommodate keychain rings. This task emphasized designing small, precise features and required the use of a vertical wall as the design context. #2 Cable management: Participants designed a holder to organize cables on a desk. The design had to securely fit the provided cables, ensuring they would not fall, and the organizer needed to remain stable when attached to the desk. #3 Toiletries holder: Participants designed a holder for a toothpaste tube and toothbrush, intended for use in a bathroom. They were provided with physical objects representing the toiletries and used the reference object functionality to ensure the holder dimensions aligned with the objects. For both conditions, participants were allowed to use a physical ruler to measure physical dimensions, although they were also allowed to use the digital AR ruler in the TinkerXR condition.
Participants engaged with TinkerXR using a Quest 3 headset and hand interactions. For TinkerXR, they were instructed to prepare and start their prints on a virtual twin of the 3D printer; while for the Tinkercad condition, they used Cura131313https://ultimaker.com/software/ultimaker-cura/’s desktop UI for 3D printing preparation. To save time, participants did not wait for the physical prints to finish before moving on to the next task. The printer slicing sequence was only carried out for the first task due to the same reason. Each modeling session was followed by a survey to gather qualitative and quantitative feedback on the tools, focusing on usability, efficiency, and overall satisfaction.
@@
Measures. We used both quantitative and qualitative metrics to evaluate the system. We recorded the time required to complete the individual 3D design tasks as a measure of overall system performance. For Task 1, we also measured the time required to transfer the finished design into the 3D printing environment and prepare it for fabrication, i.e., slicing. We measured this duration until the participant hit the print button. After completing the tasks, participants were asked to complete standardized questionnaires, including the System Usability Scale (SUS) (Brooke, 1996), which we extended with additional questions related to the immersive experience, and the NASA Task Load Index (NASA-TLX) (Hart and Staveland, 1988). We collected qualitative feedback through open-ended questions regarding their overall experience.
@@ 6.2.1. Observations and Analysis of the Fabricated Artifacts
Figure 15 shows the three types of objects designed and fabricated by the participants. Each design task had specific requirements to ensure functionality and usability, and all participants successfully met these requirements. For the key hanger, the design required at least two hooks to hold keychains. The hooks had to be narrow enough to accommodate standard keychain rings. For the cable management, participants needed to design a holder that securely fit at least two cables of provided diameters, ensuring the cables would not slip out. Additionally, the design had to accommodate the thickness of the desk to remain stable when attached. For the toiletries holder, the design required at least two compartments or holes to hold the provided toothbrush and toothpaste tube. These compartments had to match the given dimensions to ensure a stable hold. All models were validated by checking if they adhered to dimensional requirements and functionality constraints defined for each task.
A notable insight was that cube primitive as a solid and cylinder geometry as a hole were mostly used objects for both the TinkerXR and Tinkercad conditions. During TinkerXR usage, snap-to-grid was heavily utilized. All of the participants used the physical ruler for the wall-mounted key hanger and cable management tasks while designing with Tinkercad, while all of them used the reference object functionality of TinkerXR for the toiletries holder while designing with TinkerXR. Participants relied heavily on depth occlusion and the highlighted projection of the objects on the grid to understand the location and depth of the objects. Seen in Figure 15, the outcomes that have taken the most time on average of two systems were the key hanger by P5 (18m 45s), the toiletiers holder by P5 (15m 31s), and the key hanger by P9 (14m 35s).
All of the resulting objects with TinkerXR vs. Tinkercad were comparable, i.e., none of the users created designs that differed deeply depending on the tool used. This verifies that TinkerXR was able to offer similar 3D CAD capabilities despite the newly introduced immersive AR environment and not being an established commercial solution.
6.2.2. Time
Across all three design tasks, we found no statistically significant difference in task completion time using TinkerXR compared to Tinkercad: Task 1 (t = 0.272, p = 0.792), Task 2 (t = 0.272, p = 0.792), Task 3 (t = -1.769, p = 0.110). This suggests that TinkerXR provides an immersive design experience without a substantial time cost. However, future work may explore this tradeoff in larger or more varied user populations, e.g., for experts as well. Fig. 16 shows the results for all tasks. When it comes to the slicing process for Task 1, the 3D printing preparation took less time when participants utilized the TinkerXR system. A paired t-test confirmed significantly shorter completion times in the TinkerXR condition for Task 1 (t = 3.747, p = 0.004).
Result of the recorded times of the participants.